A rectangular garden is 7 yards wide. If you walked diagonally across the garden, you would walk 25 yards. How mahy yards long is the garden?
Thank you for your help.

Answer by stanbon(75887)

A rectangular garden is 7 yards wide. If you walked diagonally across the garden, you would walk 25 yards. How mahy yards long is the garden?
-----
width = 7yds
length = x yds
diagonal = 25 yds
-----
Solve for "x" using Pythagoras:
7^2 + x^2 = 25*2
-----
x^2 = 625 - 49
x = sqrt(576)
x = 24 yds (length)
=============================
Cheers,
Stan H.